# Provably Overwhelming Transformer Models with Designed Inputs

### v1

@@ D.3 Proofs and Subalgorithms for Attention Bounds
$\max_{X,X^{\prime},B,B^{\prime},S,S^{\prime}}\bigg{\|}\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{free}}\cdot(X^{\prime}[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V\bigg{\|}_{\infty}$
$\leq 2\cdot\max_{X,B,S}\bigg{\|}\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V\bigg{\|}_{\infty}$ (by triangle inequality)
$\leq 2\cdot\max_{B,S}\sum_{i\in(n_{fix},{n_{ctx}})}\vec{p}(X)_{\text{free}}[i]\cdot\mathopen{}\mathclose{{}\left\|(X[i,:]\text{E}\cdot\mathrm{diag}(B)+S)V}\right\|_{\infty}$ (by triangle inequality)
$\leq 2\cdot\max_{B,S}\max_{i}{\beta}_{free}^{\max}\cdot\mathopen{}\mathclose{{}\left\|(X[i,:]\text{E}\cdot\mathrm{diag}(B)+S)V}\right\|_{\infty}$ (by definition of $\vec{p}(X)_{\text{free}}$)
where the last inequality follows from the fact that $\vec{p}(X)_{\text{free}}$ is non-negative and the sum is at most ${\beta}_{free}^{\max}$. Finally, note that $\max_{i}{\beta}_{free}^{\max}\|X[i,:]\text{E}\cdot\mathrm{diag}(B)+S)V\|\leq{\beta}_{free}^{\max}\max_{t\in[{d_{vocab}}]}\|\text{E}[t,:]\cdot(\mathrm{diag}(B)+S)V\|$ as the maximum free token contribution is at most the maximum contribution from any possible token. Finally, for the free tokens, we get a bound
$2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S,t\in[{d_{vocab}}]}\norm{(\text{E}[t]\cdot\mathrm{diag}(B)+S)V}_{\infty}=2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(\text{E}\cdot\mathrm{diag}(B)+S)V}_{F_{\infty}}.$
@@
$\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$
@@
$\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B^{\prime})$
$+S^{\prime})V\bigg{\|}=\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\bigg{\|}(\vec{p}(X)_{\text{fix}}\cdot(\text{E}[s\cup\{q\},:]\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(\text{E}[s\cup\{q\},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$
$\leq\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$ (by the triangle inequality)
where $t_{i}$ is the $i$-th fixed token. Now, we want to re-write the last line to factor out the differing blowup and shift values.
$\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$
$=\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V$
-- $\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot(\mathrm{diag}(B)+S-(\mathrm{diag}(B)+S-\mathrm{diag}(B^{\prime})-S^{\prime}))V\bigg{\|}$
$\leq\max_{X,X^{\prime},(B,S)}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
$+$ $\max_{X,(B,S),(B^{\prime},S^{\prime})}\sum_{i}\vec{p}(X)_{\text{fix}}[i]\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-(\text{E}[t_{i},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$ (By the triangle inequality)
We now bound the two terms separately. For the second term, we have that
$\max_{X^{\prime},(B,S),(B^{\prime},S^{\prime})}\sum_{i}\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-(\text{E}[t_{i},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V\bigg{\|}$
$\leq\max_{(B,S),(B^{\prime},S^{\prime})}\max_{i}{\beta}_{fix}^{\max}\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)-(\text{E}[t_{i},:]\cdot\mathrm{diag}(B^{\prime})+S^{\prime})\bigg{\|}\cdot\norm{V}_{\infty}$ (By definition of $\vec{p}(X^{\prime})_{\text{fix}}$)
$\leq\norm{V}_{\infty}\cdot{\beta}_{fix}^{\max}\max_{t\in s\cup\{q\}}\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S}).$
For the first term, we get
$\max_{(B,S),X,X^{\prime}}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
$=\max_{(B,S),X,X^{\prime}}\sum_{i}\bigg{\|}\mathopen{}\mathclose{{}\left(\vec{p}(X)_{\text{fix}}[i]-\vec{p}(X^{\prime})_{\text{fix}}[i]}\right)\cdot(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
$=\max_{(B,S),X,X^{\prime}}\sum_{i}\mathopen{}\mathclose{{}\left(\vec{p}(X)_{\text{fix}}[i]-\vec{p}(X^{\prime})_{\text{fix}}[i]}\right)\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
$\leq\max_{(B,S),i}({\beta}_{free}^{\max}-{\beta}_{free}^{\min})\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
@@
$\max_{(B,S),i}({\beta}_{free}^{\max}-{\beta}_{free}^{\min})\cdot\bigg{\|}(\text{E}[t_{i},:]\cdot\mathrm{diag}(B)+S)V\bigg{\|}=({\beta}_{free}^{\max}-{\beta}_{free}^{\min})\cdot\max_{B,S}\norm{(\text{E}[s\cup\{q\},:]\cdot\mathrm{diag}(B)+S)V}_{F_{\infty}}.$
@@
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $\leq({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\max_{B,S}\norm{\text{E}[{s}\cup\{q\},:]\cdot\mathrm{diag}(B)+S}_{{F_{\infty}}}$
$+2{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(E\cdot\mathrm{diag}(B)+S)\cdot V}_{{F_{\infty}}}.+\norm{V}_{\infty}{\beta}_{fix}^{\max}\max_{t\in s\cup\{q\}}\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S}).$
D.4 Proofs for MLP Bounds
Proof of lemma 4.14.
The second statement follows from $\vec{e}_{n_{ctx}}\cdot f^{\mathbb{I}}_{n_{fix}\mid r,q}=f^{enc}$ directly. Then, the first statement follows from the Lipschitz constant of the MLP:
$\norm{\text{MLP}(\vec{e}_{n_{ctx}}\text{E}B+S)-\text{MLP}(\vec{e}_{n_{ctx}}\text{E}B^{\prime}+S^{\prime})}_{\infty}$ $\leq\mathrm{Lip}({\text{MLP}})_{\infty}\cdot\norm{\vec{e}_{n_{ctx}}\text{E}B+S-\vec{e}_{n_{ctx}}\text{E}B^{\prime}+S^{\prime}}_{\infty}$
$\leq\mathrm{Lip}({\text{MLP}})_{\infty}\cdot\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}.$
D.5 Proof of theorem 4.5
@@ D.4 Proofs for MLP Bounds
We first make use of theorem 4.4 to prove the first statement of the above theorem. Now, we just need to prove that the bound, $W$ in algorithm 1, is a valid bound on the lift $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$. By the monotonicity of lifting, we then have that $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}\leq\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}.$
@@
$\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $\leq\mathrm{Lip}(\mathrm{Unembed})_{\infty}\cdot(\mathcal{W}(f^{\texttt{AttnH}};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$+\mathcal{W}(f^{\text{MLP}};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}+\mathcal{W}(f^{\mathbb{I}};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty})$
@@
$f^{{\texttt{comp}}}(X,(B,S))={\texttt{comp}}\circ f^{enc}(X,(B,S))$
@@
$f^{enc}(X,(B,S))=(X\cdot\text{E}\cdot\mathrm{diag}(B)+S).$
Then, we can use lemma 4.12 to get that $W^{\texttt{Attn}}$ is a valid bound on the worst-case deviation of $f^{\texttt{AttnH}}$. Then, we use lemma 4.14 to get that $\mathrm{Lip}(\text{MLP})_{\infty}\cdot\mathcal{W}(f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})$ is a valid bound on the worst-case deviation of $f^{\text{MLP}}$. Finally, we use lemma 4.15 to get the bound on $f^{enc}$. ∎
@@ E.1 Proof of Attention Worst-Case Deviation
First note that the blowup and shift are fixed and thus $f^{enc}$ is only a function of the its input token: i.e. we can rewrite $f^{enc}(\vec{e}_{t},BS)$ as $f^{enc}(\vec{e})=\vec{e}\cdot E\cdot\mathrm{diag}(B)+S$ for fixed $B$ and $S$. Then, for the naive algorithm, algorithm 6, the correctness follows from the fact that the minimum possible logit value from each position can be calculated by enumerating over all the free tokens.
@@
Worst-case deviation of attention is bounded as follows for fixed blowup $B$ and shift $S$ when considering the permutation class $[X]$ with free tokens FreeToks:
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$\leq 2\cdot({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\norm{\text{E}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot B+S}_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\norm{(E[(n_{fix},{n_{ctx}}),:]B+S)\cdot V}_{{F_{\infty}}}.$
@@
The proof follows analogously to the proof of lemma 4.12 (bound on worst-case deviation of attention in the generic case). The main difference is that we have a singleton set for the blowup and shift and that we restrict the free tokens. First, this implies that the restriction of $\mathcal{B},\mathcal{S}$ to singelton sets means that no optimization over blowup and shift are needed in the bounds as the blowup and shift become constants. Second, the restriction of the free tokens means that the worst-case contribution from the value function, $(E\mathrm{diag}(B)+S\cdot V)$, is restricted to the embeddings selected by the free tokens. So, we replace $\norm{E\mathrm{diag}(B)+S\cdot V}_{{F_{\infty}}}$ with $\norm{(E[(n_{fix},{n_{ctx}}),:]\mathrm{diag}(B)+S)\cdot V}_{{F_{\infty}}}$. Moreover, $\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S})_{\infty}=0$ as the blowup and shift are fixed and so we can drop the last term in the bound in lemma 4.12. ∎
@@
then the output of model $\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q}$ is fixed for all inputs in $[X]$. Moreover, we algorithm 2 in appendix E produces an upper bound $W$ for $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$.
@@
The proof follows analogously to the proof of theorem 4.5 in section D.5. The key difference is that, due to the fixing of the blowup and shift, the bound on $W$ simplifies. First, the worst-case deviation of $\vec{e}_{{n_{ctx}}}\cdot f^{enc}(X)$ is 00 as the blowup and shift are fixed. So, the worst-case deviation of the MLP and identity function are 00 as well. Then, algorithm uses either algorithm 6 or algorithm 3 to find bounds on the extremal values of attention with provable correctness as per lemma 5.3. Next, the algorithm simplified worst-case deviation of attention (as per lemma 5.4) to bound the worst-case deviation of attention. Finally, the algorithm uses the Lipschitz constant of the unembedding function, invoking lemma C.1, to bound the worst-case deviation of the model. ∎

### v2

@@ D.1 Proofs and Subalgorithms for Attention Bounds
$\max_{X,X^{\prime}}\bigg{\|}\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}_{\texttt{LN}})V-\vec{p}(X^{\prime})_{\text{free}}\cdot(X^{\prime}[(n_{fix},{n_{ctx}}):,]\text{E}_{\texttt{LN}})V\bigg{\|}_{\infty}$
$\leq 2\cdot\max_{X}\bigg{\|}\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}_{\texttt{LN}})V\bigg{\|}_{\infty}$ (by triangle inequality)
$\leq 2\sum_{i\in(n_{fix},{n_{ctx}})}\vec{p}(X)_{\text{free}}[i]\cdot\mathopen{}\mathclose{{}\left\|(X[i,:]\text{E}_{\texttt{LN}})V}\right\|_{\infty}$ (by triangle inequality)
$\leq 2\cdot\max_{i}{\beta}_{free}^{\max}\cdot\mathopen{}\mathclose{{}\left\|(X[i,:]\text{E}_{\texttt{LN}})V}\right\|_{\infty}$ (by definition of $\vec{p}(X)_{\text{free}}$)
where the last inequality follows from the fact that $\vec{p}(X)_{\text{free}}$ is non-negative and the sum is at most ${\beta}_{free}^{\max}$. Finally, note that $\max_{i}{\beta}_{free}^{\max}\|X[i,:]\text{E}_{\texttt{LN}})V\|\leq{\beta}_{free}^{\max}\max_{t\in[{d_{vocab}}]}\|\text{E}[t,:]\cdot V\|$ as the maximum free token contribution is at most the maximum contribution from any possible token. Finally, for the free tokens, we get a bound
$2\cdot{\beta}_{free}^{\max}\cdot\max_{t\in[{d_{vocab}}]}\norm{(\text{E}_{\texttt{LN}}[t])V}_{\infty}=2\cdot{\beta}_{free}^{\max}\cdot\norm{(\text{E}_{\texttt{LN}})V}_{F_{\infty}}.$
@@
$\max_{X,X^{\prime}}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}_{\texttt{LN}})V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}_{\texttt{LN}})V\bigg{\|}$
@@
$\max_{X,X^{\prime}}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}_{\texttt{LN}})V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}_{\texttt{LN}}\cdot)V\bigg{\|}$
$=\max_{X,X^{\prime}}\bigg{\|}(\vec{p}(X)_{\text{fix}}\cdot(\text{E}_{\texttt{LN}}[s\cup\{q\},:])V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(\text{E}_{\texttt{LN}}[s\cup\{q\},:])V\bigg{\|}$
$\leq\max_{X,X^{\prime}}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}_{\texttt{LN}}[t_{i},:])V-\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}_{\texttt{LN}}[t_{i},:])V\bigg{\|}$ (by the triangle inequality)
where $t_{i}$ is the $i$-th fixed token. And then, we get
$\max_{X,X^{\prime}}\sum_{i}\bigg{\|}\vec{p}(X)_{\text{fix}}[i]\cdot(\text{E}_{\texttt{LN}}[t_{i},:])V-\vec{p}(X^{\prime})_{\text{fix}}[i]\cdot(\text{E}_{\texttt{LN}}[t_{i},:])V\bigg{\|}$
$=\max_{X,X^{\prime}}\sum_{i}\bigg{\|}\mathopen{}\mathclose{{}\left(\vec{p}(X)_{\text{fix}}[i]-\vec{p}(X^{\prime})_{\text{fix}}[i]}\right)\cdot(\text{E}_{\texttt{LN}}[t_{i},:])V\bigg{\|}$
$=\max_{X,X^{\prime}}\sum_{i}\mathopen{}\mathclose{{}\left(\vec{p}(X)_{\text{fix}}[i]-\vec{p}(X^{\prime})_{\text{fix}}[i]}\right)\cdot\bigg{\|}(\text{E}_{\texttt{LN}}[t_{i},:])V\bigg{\|}$
$\leq\max_{i}({\beta}_{free}^{\max}-{\beta}_{free}^{\min})\cdot\bigg{\|}(\text{E}_{\texttt{LN}}[t_{i},:])V\bigg{\|}$
@@
$\max_{i}({\beta}_{free}^{\max}-{\beta}_{free}^{\min})\cdot\bigg{\|}(\text{E}[t_{i},:])V\bigg{\|}=({\beta}_{free}^{\max}-{\beta}_{free}^{\min})\cdot\max_{B,S}\norm{(\text{E}[s\cup\{q\},:])V}_{F_{\infty}}.$
@@
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $\leq({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\max_{B,S}\norm{\text{E}[{s}\cup\{q\},:]\cdot V}_{{F_{\infty}}}$
$+2{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(\text{E}_{\texttt{LN}})\cdot V}_{{F_{\infty}}}.$
D.2 Proof of theorem 4.5
@@
We first make use of theorem 4.4 to prove the first statement of the above theorem. Now, we just need to prove that the bound, $W$ in algorithm 1, is a valid bound on the lift $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}$.
@@
$\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}$ $\leq\mathrm{Lip}(\mathrm{Unembed})_{\infty}\cdot(\mathcal{W}(f^{\texttt{AttnH}};\mathbb{O}_{des})_{\infty}$
$+\mathcal{W}(f^{\text{MLP}};\mathbb{O}_{des})_{\infty}+\mathcal{W}(f^{\mathbb{I}};\mathbb{O}_{des})_{\infty})$
@@
$f^{{\texttt{comp}}}(X)={\texttt{comp}}\circ f^{enc}(X)$
@@
$f^{enc}(X)=(X\cdot\text{E}_{\texttt{LN}}).$
Then, we can use lemma 4.11 to get that $W^{\texttt{Attn}}$ is a valid bound on the worst-case deviation of $f^{\texttt{AttnH}}$. Then, we use lemma 4.12 to get that $\mathrm{Lip}(\text{MLP})_{\infty}\cdot\mathcal{W}(f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})$ is a valid bound on the worst-case deviation of $f^{\text{MLP}}$. ∎
@@ E.1 Proof of Attention Worst-Case Deviation
First, note that algorithm 5 is the same as algorithm 4 except that we restrict our free-token attention to those specified by the permutation class $[X]$. Thus, the correctness of algorithm 4 implies that algorithm 5 is correct.
@@
Worst-case deviation of attention is bounded as follows when considering the permutation class $[X]$ with free tokens freeToks:
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des})_{\infty}$
$\leq({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\norm{\text{E}_{\texttt{LN}}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot V\cdot\mathrm{Unembed}}_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\norm{(\text{E}_{\texttt{LN}}[(n_{fix},{n_{ctx}}),:]\cdot V\cdot\mathrm{Unembed}}_{{F_{\infty}}}.$
@@
Again, for notational simplicity, we will absorb the $\mathrm{Unembed}$ matrix into the value matrix $V$. We will write $V=V^{\prime}\cdot\mathrm{Unembed}$ where $V^{\prime}$ is the value matrix in the attention head. The proof is the same as the proof of lemma 4.11 (bound on worst-case deviation of attention in the generic case). ∎
@@
then the output of model $\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q}$ is fixed for all inputs in $[X]$. Moreover, we algorithm 2 in appendix E produces an upper bound $W$ for $\mathcal{W}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}$.
@@
The proof follows analogously to the proof of theorem 4.5 in section D.2 except that we consider the permutation class $[X]$ and thus can use either the algorithms algorithm 5 or algorithm 3 to find bounds on ${\alpha}_{free}^{\min}$ and ${\alpha}_{free}^{\max}$. ∎